OpenStudy (anonymous):

Use the distributive property to simplify 8(3x + 1).

OpenStudy (anonymous):

you need \[8\times 3x+8\times 1\]

OpenStudy (anonymous):

8(3x)+8(1) = 24x+8

OpenStudy (aravindg):

24xx+8